High-Low

Biblio Files Company is the chief competitor of Cover-to-Cover Company in the bookshelf business. Biblio Files is analyzing its manufacturing costs, and has compiled the following data for the first six months of the year. After reviewing the data, answer questions (1) through (3) that follow.

Units Produced Total Cost
January 4,360 units $65,600
February 300 6,250
March 1,000 15,000
April 5,800 116,250
May 1,750 32,500
June 3,015 48,000

1. From the data previously provided, help Biblio Files Company estimate the fixed and variable portions of its total costs using the high-low method. Recall that Total Costs = (Variable Cost Per Unit x Number of Units Produced) + Fixed Cost. Complete the following table.

Total Fixed Cost Variable Cost per Unit
$ $
Answer #1

Variable cost per unit = (Highest activity cost - Lowest activity cost) / (Highest activity unit - Lowest activity unit)

= (116,250-6,250) / ( 5,800-300)

= 110,000/5,500

= $20 per unit

Fixed cost = Highest activity cost - ( Variable cost per unit x Highest activity unit)

= 116,250 - (20 x 5,800)

= 116,250-116,000

= $250

Total Fixed Cost Variable Cost per Unit
$250 $20

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks